Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2023 March Paper 52 Q2

A Level / CIE / S1

CIE 9709 2023 March Paper 52 Paper · Question 2

题目

Problem

Alisha has four coins. One of these coins is biased so that the probability of obtaining a head is 0.6. The other three coins are fair. Alisha throws the four coins at the same time. The random variable XX denotes the number of heads obtained.

(a) Show that the probability of obtaining exactly one head is 0.225.

(3)

(b) Complete the following probability distribution table for XX.

(2)

xx0011223344
P(X=x)\mathrm{P}(X=x)0.050.050.2250.2250.0750.075

(c) Given that E(X)=2.1\mathrm{E}(X)=2.1, find the value of Var⁡(X)\operatorname{Var}(X).

(2)

题目中文翻译

Alisha 有四枚硬币。其中一枚有偏,出现正面的概率为 0.6,另外三枚是均匀硬币。Alisha 同时抛掷这四枚硬币。随机变量 XX 表示得到的正面数。

(a) 证明恰好得到一个正面的概率为 0.225。

(3)

(b) 补全下面 XX 的概率分布表。

(2)

xx0011223344
P(X=x)\mathrm{P}(X=x)0.050.050.2250.2250.0750.075

(c) 已知 E(X)=2.1\mathrm{E}(X)=2.1,求 Var⁡(X)\operatorname{Var}(X) 的值。

(2)

原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

按偏币是正面还是反面分类:偏币正则其余三枚全反,偏币反则其余三枚恰一正。

答题过程

展开

If the biased coin is a head, all three fair coins must be tails, with probability 0.6(0.5)30.6(0.5)^3.

If the biased coin is a tail, exactly one fair coin must be a head, with probability 0.4(31)(0.5)30.4\binom31(0.5)^3.

Thus

P(X=1)=0.6(0.5)3+0.4(3)(0.5)3=0.075+0.150=0.225.\mathrm{P}(X=1)=0.6(0.5)^3+0.4(3)(0.5)^3 =0.075+0.150=\boxed{0.225}.

(b)

解法一

思路

展开

同样按偏币结果分类,分别计算总正面数为2与3的概率。

答题过程

展开 P(X=2)= 0.6(31)(0.5)3+0.4(32)(0.5)3= 0.225+0.150=0.375,\begin{align*} \mathrm{P}(X=2) =&\,0.6\binom31(0.5)^3+0.4\binom32(0.5)^3\\[4mm] =&\,0.225+0.150=0.375, \end{align*} P(X=3)= 0.6(32)(0.5)3+0.4(0.5)3= 0.225+0.050=0.275.\begin{align*} \mathrm{P}(X=3) =&\,0.6\binom32(0.5)^3+0.4(0.5)^3\\[4mm] =&\,0.225+0.050=0.275. \end{align*}

The completed distribution is

xx0011223344
P(X=x)\mathrm{P}(X=x)0.050.050.2250.2250.3750.3750.2750.2750.0750.075

(c)

解法一

思路

展开

先算二阶矩,均值直接用题给2.1。

答题过程

展开 E(X2)= 0.225+4(0.375) +9(0.275)+16(0.075)= 5.4.\begin{align*} \mathrm{E}(X^2)=&\,0.225+4(0.375)\\[4mm] &\,\hspace{2pt}+9(0.275)+16(0.075)\\[4mm] =&\,5.4. \end{align*}

Hence

Var⁡(X)=5.4−(2.1)2=0.99.\operatorname{Var}(X)=5.4-(2.1)^2=\boxed{0.99}.